OpenStudy (anonymous):

2x-3=-24 and 3x+2y=-8

OpenStudy (anonymous):

Solve for x first then substitute the answer in the second equation to get the value of y. 2x-3=-24 2x=-21 x=-10.5 3(-10.5)+2y=-8 -31.5+2y=-8 2y=23.5 y=11.75

OpenStudy (anonymous):

Is it parallel, perpendicular or neither

OpenStudy (anonymous):

Wait-was there supposed to be a y in the first equation?

OpenStudy (anonymous):

Yes, it's determining for slope

OpenStudy (anonymous):

Disregard what I said earlier. Can you include the first equation with y?

OpenStudy (anonymous):

Can you rewrite the first equation with y. I need to no if it was after the -3 or -24.

OpenStudy (anonymous):

know*

OpenStudy (anonymous):

I'm going to solve the second equation in terms of y. 3x+2y=-8 2y=-3x-8 y=-3/2-4 You said there was a y in the first equation, but you didn't tell me where it was. I'm going to assume it was after the -3. 2x-3y=-24 -3y=-2x-24 y=2/3x+8 In this case, they would be perpendicular, because their slopes are negative reciprocals of each other. (If they were the same, they would be parallel and if they were completely different, they would be neither.) I hope I helped.

OpenStudy (anonymous):

Thank you... Sorry I didn't get back. I'm at work

OpenStudy (anonymous):

That's fine. Was the y after the -3?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay. So it's perpendicular. Have a good day. :)

OpenStudy (anonymous):

Thanks you too